OpenStudy (anonymous):

is it A? Find the product of (x + 3)(x − 3). x2 − 6x + 9 x2 + 6x + 9 x2 + 9 x2 − 9

jimthompson5910 (jim_thompson5910):

use the difference of squares rule

OpenStudy (anonymous):

I have a problem with the adding and subtracting

jimthompson5910 (jim_thompson5910):

the difference of squares rule is (a+b)(a-b) = a^2 - b^2

jimthompson5910 (jim_thompson5910):

does that help?

OpenStudy (anonymous):

yess thank you! so its A?

jimthompson5910 (jim_thompson5910):

no it's not A,use the template I gave you

OpenStudy (anonymous):

its B

jimthompson5910 (jim_thompson5910):

similar example (x-5)(x+5) = x^2 - 5^2 (x-5)(x+5) = x^2 - 25

OpenStudy (texaschic101):

do you know the FOIL method ?

OpenStudy (anonymous):

yeah I did it its B

jimthompson5910 (jim_thompson5910):

it's not B either

OpenStudy (texaschic101):

your just guessing....work it out. Do you know the FOIL method or you can just do it the way jim showed you

OpenStudy (texaschic101):

do you want some help ?

OpenStudy (texaschic101):

I am not going to participate unless you participate

OpenStudy (anonymous):

yes please, I will

jimthompson5910 (jim_thompson5910):

do you see how my example works ?

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

so in the case of the example a = x b = 5 and they are plugged into (a+b)(a-b) = a^2 - b^2

OpenStudy (anonymous):

so x^2-5^2

jimthompson5910 (jim_thompson5910):

correct, then 5^2 turns into 25

jimthompson5910 (jim_thompson5910):

so that's how I got (x-5)(x+5) = x^2 - 25

OpenStudy (texaschic101):

jim...this doesn't work on all problems, should I show her the FOIL method or do you thin that will confuse her

OpenStudy (anonymous):

so x^2+10x+25

OpenStudy (texaschic101):

typo * think

jimthompson5910 (jim_thompson5910):

go ahead I figured that the short and easy way works just as well

OpenStudy (anonymous):

x^2+10+25x?

jimthompson5910 (jim_thompson5910):

but you can go the FOIL route if you want

OpenStudy (texaschic101):

ok.....I like this way FOIL it stands for (first, outside, inside, last. (x + 3)(x - 3) F - multiply the first terms in each set....x * x = x^2 you can do the rest O - multiply the outside terms in each set...x * -3 = ? I - multiply the inside terms in each set...3 * x = ? L - multiply the last terms in each set...3 * -3 = ? Can you understand what I am doing ?

OpenStudy (anonymous):

yes but brb I gotta go to the store real quick

OpenStudy (texaschic101):

tag me when you get back and we can finish

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

I am here....shall we finish...can you answer the last questions I posted ?

OpenStudy (anonymous):

x^2-9x-9

OpenStudy (texaschic101):

not quite...don't guess....lets do this we are on the outside terms O - multiply the outside terms in each set...x * -3 = ?

OpenStudy (anonymous):

-3x

OpenStudy (texaschic101):

correct....now the inside terms...3 * x = ?

OpenStudy (anonymous):

3x

OpenStudy (texaschic101):

correct....now the last terms in each set...3 * -3 = ?

OpenStudy (anonymous):

-9

OpenStudy (texaschic101):

Now we will put all our answers together... x^2 - 3x + 3x - 9 now we can combine like terms can you do that ?

OpenStudy (anonymous):

yes so its x^2+9x-9

OpenStudy (texaschic101):

no....look at your -3x and 3x....they are opposites so they cancel each other out leaving us with x^2 - 9 Do you understand ?

OpenStudy (anonymous):

ohhhh, yeah now I understand lol

OpenStudy (texaschic101):

thats good....Just remember FOIL....first,outside,inside,last....it will come easier once you learn it....

OpenStudy (anonymous):

I think I'm getting it now, thank you so much!

OpenStudy (texaschic101):

your very welcome...glad I could help :)